OpenStudy (anonymous):

Sharita does special typing jobs for scientists. She charges $7 for the first 3 pages and $3 for each additional page of scientific typing. Sonia, a biologist, paid $97 to have a research paper typed. How many pages long was Sonia's paper?

OpenStudy (mathmusician):

Let x be the number of pages: 100 = 7 + 3(x - 3) 100 = 7 + 3x - 9 102 = 3x x = 102/3 = 34 pages check: 7 + 3(31) = 100 checks The paper was 34 pages.

OpenStudy (anonymous):

thanks but thats wrong

OpenStudy (anonymous):

The First 3 pages cost $7 and the remaining pages cost $3 each, for a total of $97. Let n = the number of pages typed and n - 3 = the number typed at $3. @MathMusician